# Multi-Token Joint Speculative Decoding for Accelerating Large Language Model Inference

### v1

@@ 3.2 Multi-Token Joint Speculative Decoding
Unfortunately, the computation cost of MJGD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$. Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” [20, 17], we propose multi-token joint speculative decoding (MJSD), which exploits a small model $M_{q}$ to accelerate MJGD approximately. The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^4] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$. Without loss of generality, we consider non-warped distribution in the illustration of MJSD. and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration. However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$. So we propose to further approximate this process with beam decoding, which is an effective and efficient algorithm to generate sequences with high likelihood. In this way, MJSD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does. Algorithm 1 in the Appendix illustrates the pseudocode of MJSD algorithm.
@@
Figure 2: An example of MJSD’s verification process. MJSD accepts the longest draft sub-sequence that passes verification based on joint likelihood.
@@
Figure 2 illustrates the verification process of MJSD. Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam decoding with the small model. Since beam decoding is a widely recognized algorithm to generate sequences with high overall likelihood [18], it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large. Also, since beam decoding works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large. To approximate MJGD, for each step $i$, MJSD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$. So MJSD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$. Then for each prefix sub-sequence $x_{t+1:t+j}$, it passes verification if and only if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, where $\tau\in[0,1)$ is a pre-defined threshold. Notice that if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, we have $\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})}>\tau$, which means $\frac{q(x_{t+1:t+j}|x_{1:t})-p(x_{t+1:t+j}|x_{1:t})}{p(x_{t+1:t+j}|x_{1:t})}<\frac{1}{\tau}-1$. Therefore, our acceptance policy guarantees that when $q(x_{t+1:t+j}|x_{1:t})>p(x_{t+1:t+j}|x_{1:t})$, the relative error is bounded. And if $q(x_{t+1:t+j}|x_{1:t})\leq p(x_{t+1:t+j}|x_{1:t})$, it means the sub-sequence has higher likelihood in the large model, then it is reasonable to accept it. After verifying all the sub-sequences, MJSD accepts the longest draft sub-sequence that passes verification.
@@
The verification step of MJSD ensures that the accepted tokens have a high joint likelihood with the large model. We have shown that selecting multiple tokens based on their joint likelihood lead to better output perplexity. Thus, MJSD is more effective than greedy decoding and vanilla speculative decoding. Furthermore, since MJSD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high. So at each iteration, MJSD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
@@ 4 Energy Efficiency Analysis
Previous studies [20, 5, 17, 32] only focus on the speed of speculative decoding. However, an equally important consideration is energy consumption. To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption. Although MJSD and speculative decoding increase the number of FLOPs due to the involvement of an additional small model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption [1, 6]. Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
To understand the net effect of speculative decoding, we decompose the total energy consumption into two parts following [1]:
@@
To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory [16]. It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b [25], requires about 1GB memory. Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs [1]. So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$. Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs. During greedy decoding and speculative decoding, the batch size is usually small [20]. So most of the computing power is not utilized [20], which means $PW_{mem}\gg PW_{flop}$.
In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ [20]. Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$. Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference. We ran OPT-13b models on 2 Nvidia A10 GPUs with 24GB memory. Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost. The details of energy measurement are illustrated in the Appendix B. Table 1 shows the results. As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly. This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$. In the next section, we will show more direct evidence for the energy efficiency of MJSD and other speculative decoding algorithms.
@@ 5 Experiments
Datasets and Models.We use three public datasets for evaluation: (1) ChatGPT-Prompt [26], (2) ChatAlpaca [3], and (3) CNN Dailymail [30]. The three datasets represent three common tasks for LLMs: instruction, multi-turn chat, and summarization. Table 4 in the Appendix shows more details of the datasets. We use two pairs of small and large models in our experiments: OPT-125m and OPT-13B [38], and Llama-68M [24] and Llama-2-13B [34].
Baselines. For each pair of small and large models, we compare our method with greedy decoding (greedy), and three speculative decoding methods: vanilla speculative decoding (speculative[19, 5], Spectr [32], and BiLD [17]. Our implementation of MJSD and all the baselines are based on a public implementation of speculative decoding [2]. We also tried to compare with SpecInfer [24] using their released implementation. But it runs out of memory on our machine. For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds. We open-sourced our code for reproduction.
More details of the hyper-parameters, warping operations, and machine configurations of the experiments can be found in the Appendix B, C, and D .
Table 2: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), and CNNDailyMail (CD) datasets. Bold numbers mark the best result, underlined numbers mark the second best.
@@
speed up $\uparrow$ 1.00$\pm$0.00 1.63$\pm$0.02 1.52$\pm$0.04 1.99$\pm$0.06 2.78$\pm$0.01
energy (J/token) $\downarrow$ 11.7$\pm$1.56 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 3.38$\pm$0.02
perplexity $\downarrow$ 3.74$\pm$0.14 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 2.06$\pm$0.06
OPT speed (token/s) $\uparrow$ 22.4$\pm$0.48 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 55.8$\pm$0.30
speed $\uparrow$ 1.00$\pm$0.00 1.51$\pm$0.08 1.41$\pm$0.06 1.70$\pm$0.06 2.50$\pm$0.05
energy (J/token) $\downarrow$ 13.2$\pm$0.28 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 3.61$\pm$0.03
perplexity $\downarrow$ 5.49$\pm$0.15 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 3.00$\pm$0.09
CA Llama-2 speed (token/s) $\uparrow$ 22.0$\pm$0.15 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 44.1$\pm$0.25
speed up $\uparrow$ 1.00$\pm$0.00 1.43$\pm$0.02 1.31$\pm$0.01 1.26$\pm$0.00 2.00$\pm$0.02
energy (J/token) $\downarrow$ 11.2$\pm$0.16 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 4.72$\pm$0.03
perplexity $\downarrow$ 2.11$\pm$0.01 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 1.88$\pm$0.05
OPT speed (token/s) $\uparrow$ 23.8$\pm$0.10 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 49.6$\pm$0.42
speed $\uparrow$ 1.00$\pm$0.00 1.49$\pm$0.01 1.62$\pm$0.03 1.19$\pm$0.01 2.08$\pm$0.03
energy (J/token) $\downarrow$ 11.3$\pm$0.22 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 4.03$\pm$0.02
perplexity $\downarrow$ 3.28$\pm$0.06 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 2.07$\pm$0.03
CD Llama-2 speed (token/s) $\uparrow$ 21.7$\pm$0.08 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 44.2$\pm$0.99
speed up $\uparrow$ 1.00$\pm$0.00 1.41$\pm$0.00 1.41$\pm$0.01 1.15$\pm$0.01 2.04$\pm$0.05
energy (J/token) $\downarrow$ 11.3$\pm$0.21 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 4.80$\pm$0.12
perplexity $\downarrow$ 2.88$\pm$0.04 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.63$\pm$0.10
OPT speed (token/s) $\uparrow$ 23.3$\pm$0.81 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 43.6$\pm$0.33
speed $\uparrow$ 1.00$\pm$0.00 1.36$\pm$0.04 1.32$\pm$0.04 1.01$\pm$0.02 1.87$\pm$0.02
energy (J/token) $\downarrow$ 11.5$\pm$0.14 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 4.86$\pm$0.03
perplexity $\downarrow$ 3.93$\pm$0.14 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.17$\pm$0.06

### v2

@@ 3.2 Multi-Token Assisted Decoding
Unfortunately, the computation cost of MTJD is infeasible in practice, since the time and space complexity to compute the joint distribution of $\gamma_{i}$ tokens is $|V|^{\gamma_{i}}$. Inspired by speculative decoding and the facts that “even when a small model is an order of magnitude smaller than a large model, only a small fraction of the small model’s prediction deviate from those of the large model” (Leviathan et al., 2023; Kim et al., 2023), we propose multi-token assisted decoding (MTAD), which exploits a small auxiliary model $M_{q}$ to accelerate MTJD approximately. The core idea is to (1) use the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$ output by $M_{q}$ to approximate $p(x_{t+1:t+\gamma_{i}}|x_{1:t})$ [^3] It is also valid to approximate $\tilde{p}$ with $\tilde{q}$. Without loss of generality, we consider non-warped distribution in the illustration of MTAD. and generate $\gamma$ draft tokens from $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$, then (2) use the large model to validate draft tokens in parallel and accept the longest draft prefix sub-sequence that passes verification, and (3) sample an additional token from the distribution of the large model without extra overhead to ensure at least one token is generated at each iteration. However, it is still infeasible to directly generate draft tokens from the joint distribution $q(x_{t+1:t+\gamma_{i}}|x_{1:t})$. So we propose to further approximate this process with beam sampling, which is an effective and efficient algorithm to generate sequences with high likelihood. In this way, MTAD reduces the number of runs of the large model to generate $N$ tokens, thus accelerating the inference in the same way as vanilla speculative decoding does. Algorithm 1 in the Appendix illustrates the pseudocode of MTAD algorithm.
@@
Figure 2: An example of MTAD’s verification process. MTAD accepts the longest draft sub-sequence that passes verification based on joint likelihood.
@@
Figure 2 illustrates the verification process of MTAD. Let $x_{t+1},\ldots,x_{t+\gamma}$ be the draft tokens generated by beam sampling with the auxiliary model. Since beam sampling is a widely recognized algorithm to generate sequences with high overall likelihood (Leblond et al., 2021), it is reasonable to assume $q(x_{t+1:t+\gamma}|x_{1:t})$ is large. Also, since beam sampling works in an autoregressive way, we can also assume that $\forall j\in\{1,\ldots,\gamma\}$, $q(x_{t+1:t+j}|x_{1:t})$ is large. To approximate MTJD, for each step $i$, MTAD needs to ensure the accepted tokens $x_{t+1:t+\gamma_{i}}$ ($0\leq\gamma_{i}\leq\gamma$) also have high joint likelihood with the large model $M_{p}$. So MTAD first computes the joint likelihood $p(x_{t+1:t+j}|x_{1:t})$ for $j=1,\ldots,\gamma$. Then for each prefix sub-sequence $x_{t+1:t+j}$, it passes verification if and only if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, where $\tau\in[0,1)$ is a pre-defined threshold. Notice that if $\min(1,\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})})>\tau$, we have $\frac{p(x_{t+1:t+j}|x_{1:t})}{q(x_{t+1:t+j}|x_{1:t})}>\tau$, which means $\frac{q(x_{t+1:t+j}|x_{1:t})-p(x_{t+1:t+j}|x_{1:t})}{p(x_{t+1:t+j}|x_{1:t})}<\frac{1}{\tau}-1$. Therefore, our acceptance policy guarantees that when $q(x_{t+1:t+j}|x_{1:t})>p(x_{t+1:t+j}|x_{1:t})$, the relative error is bounded. And if $q(x_{t+1:t+j}|x_{1:t})\leq p(x_{t+1:t+j}|x_{1:t})$, it means the sub-sequence has higher likelihood in the large model, then it is reasonable to accept it. After verifying all the sub-sequences, MTAD accepts the longest prefix sub-sequence that passes verification.
@@
The verification step of MTAD ensures that the accepted tokens have a high joint likelihood with the large model. We have shown that selecting multiple tokens based on their joint likelihood lead to better output perplexity. Thus, MTAD is more effective than multinomial sampling and vanilla speculative decoding. Furthermore, since MTAD accepts the longest draft sub-sequence with high likelihood, it can tolerate low-quality tokens as long as the joint likelihood is high. So at each iteration, MTAD can accept more draft tokens than vanilla speculative decoding, which results in better efficiency.
Next, we theoretically analyze the approximation error of MTAD. Lemma 3.4 shows the upper bound of MTAD’s perplexity. And Theorem 3.5 shows the upper bound of the ratio between the perplexity of approximate MTAD and exact MTJD. The proofs are given in Appendix A.
Lemma 3.4.
Let us assume that when the small auxiliary model generates draft tokens with beam sampling, the beam width is large enough such that the returned log-likelihood is close to the maximum log-likelihood, i.e.,
$\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))\geq(1-\epsilon)\mathbb{E}_{x_{1:\Gamma_{i-1}}\in\mathcal{X}}\max_{x_{\Gamma_{i-1}+1:\Gamma_{i}-1}}\log q(x_{\Gamma_{i-1}+1:\Gamma_{i}-1}|x_{1:\Gamma_{i-1}}))$ (5)
where $\epsilon$ is an error term and $\epsilon\leq 0$ because $\log q\leq 0$.
Furthermore, let $H(p,q)$ the single-token cross entropy between $p$ and $q$, i.e., $H(p,q)=-\mathbb{E}_{x_{1:t\in\mathcal{X}}}\sum_{x_{t+1}}p(x_{t+1}|x_{1:t})\log q(x_{t+1}|x_{1:t})$.
With the two assumption above, when $N\rightarrow\infty$ we have
$PPL_{q}({x_{1:\Gamma_{N}}})\leq\exp(-\frac{1-\epsilon}{\bar{\gamma}}\mathbb{E}_{\gamma}L_{q}(\gamma-1,\arg\max\circ q)+\frac{H(p,q)}{\bar{\gamma}})$ (6)
where
$L_{q}(\gamma,\arg\max\circ q)=\mathbb{E}_{x_{1:t}\in\mathcal{X}}\max_{x_{t+1:t+\gamma}}\log q(x_{t+1:t+\gamma}|x_{1:t}))$ (7)
Theorem 3.5.
Let $x_{1:\Gamma_{N}}$ be the tokens generated by approximate MTAD, and $x^{*}_{1:\Gamma_{N}}$ be the tokens generated by deterministic exact MTJD. Assume $\forall x_{1:t}\in\mathcal{X}$, $\|\log p(x|x_{1:t})-\log q(x|x_{1:t})\|_{\infty}\leq U$, where $U$ is a constant. We have
$\lim_{N\rightarrow\infty}\frac{PPL_{p}(x_{1:\Gamma_{N}})}{PPL_{p}(x^{*}_{1:\Gamma_{N}})}\leq\tau^{-\frac{1}{\bar{\gamma}}}\exp\left(\frac{(1-\epsilon\bar{\gamma})H(p)+(1-\epsilon+\bar{\gamma})U}{\bar{\gamma}}\right)$ (8)
where $H(p)$ is the entropy of $p$ and $\epsilon<0$ is the error term of beam sampling (see Lemma 3.4).
Theorem 3.5 suggests the approximation error of MTAD is bounded by a factor related to the verification threshold $\tau$, average number of accepted tokens $\bar{\gamma}$, the difference between the large and small models (measured by $U$), the error of beam sampling $\epsilon$, and the entropy of the large model itself. In addition, the following theorem analyzes $\bar{\gamma}$. The proof is illustrated in Appendix A.
Theorem 3.6.
Following the assumption in Theorem 3.5, we have $\bar{\gamma}\geq\frac{|\log\tau|}{U}$.
With Theorem 3.6, we observe that when $q\rightarrow p$, we have $U\rightarrow 0$ and $\bar{\gamma}\rightarrow\infty$. Meanwhile, when $\epsilon\rightarrow 0$, meaning the beam width for the auxiliary model is large enough, the ratio bound in Theorem 3.5 converges to 1, It implies that MTAD converges to MTJD under these limiting conditions.
Similar to Spectr (Sun et al., 2023) and SpecInfer (Miao et al., 2023), it is possible to enhance the number of accepted tokens in MTAD by allowing the draft model to generate multiple draft sequences and applying tree-based attention (Miao et al., 2023) for simultaneous verification. However, our preliminary experiment results suggest that since MTAD already selects the longest accepted prefix sub-sequence, the advantage of generating multiple draft tokens is less significant. Moreover, this approach increases the memory cost during inference and may affect the error bounds derived above. Therefore, we leave a more detailed exploration of this extension as future work.
@@ 4 Energy Efficiency Analysis
Previous studies (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023; Sun et al., 2023) only focus on the speed of speculative decoding. However, an equally important consideration is energy consumption. To our knowledge, there is no existing work evaluating the impact of speculative decoding on inference energy consumption. Although MTAD and speculative decoding increase the number of FLOPs due to the involvement of a small auxiliary model and the rollback operation, they concurrently reduce the inference time and memory operations, which are key factors of GPU energy consumption (Allen & Ge, 2016; Chen et al., 2011). Consequently, it poses an open question regarding whether speculative decoding increases or decreases overall energy consumption.
To understand the net effect of speculative decoding, we decompose the total energy consumption into two parts following (Allen & Ge, 2016):
@@
Table 3: Speed and energy cost of multinomial sampling (ms) and speculative decoding (spec).
OPT Llama-2
ms spec ms Spec
tokens/s 23.8 35.6 22.0 31.6
J/token 11.3 5.74 11.2 6.97
To determine the relative magnitude relationship between $PW_{flop}$ and $PW_{mem}$, we first point out the fact that GPU memory operations in LLM inference are dominated by accessing off-chip global memory, which consumes about 100×100$\times$100 × of energy compared to accessing on-chip shared memory (Jouppi et al., 2021). It is because each multiprocessor on GPU usually has 64KB of on-chip memory shared by multiple threads, while storing a single layer of LLM, say T5-11b (Raffel et al., 2020), requires about 1GB memory. Moreover, Allen and Ge showed that doing sequential read from off-chip memory consumes $20$-$30\%$ more power than running maximum FLOPs (Allen & Ge, 2016). So we have $PW_{mem}>PW^{*}_{flop}\geq PW_{flop}$. Notice that $PW^{*}_{flop}=PW_{flop}$ only if the batch size reaches the maximum parallelization capacity of GPUs. During multinomial sampling and speculative decoding, the batch size is usually small (Leviathan et al., 2023). So most of the computing power is not utilized (Leviathan et al., 2023), which means $PW_{mem}\gg PW_{flop}$.
In addition, previous studies have shown that during LLM inference $T_{mem}\gg T_{flop}$ (Leviathan et al., 2023). Therefore, the energy induced by memory operations, i.e., $PW_{mem}T_{mem}$ dominates $E_{total}$. Since speculative decoding reduces $T_{mem}$ by reducing the number of runs of the large model, it should reduce the inference energy consumption to a similar extent as it reduces time consumption.
To validate our hypothesis, we conducted an experiment to evaluate how batch size influences energy consumption during inference. We ran OPT-13b models on a Nvidia L40 GPUs with 48GB memory. Fixing the total number of runs of the large model while varying the input batch size $b\in\{1,2,4,8,16\}$ for each run, we measured time and energy cost. The details of energy measurement are illustrated in the Appendix D. Table 2 shows the results. As batch size doubles, although the number of FLOPs doubles, the energy consumption per run increases slightly. This observation demonstrates that $PW_{mem}T_{mem}$ dominates $E_{total}$. Moreover, we measured the speed and energy consumption of running multinomial sampling with the large model and speculative decoding using OPT (125M, 13B) and Llama-2 (68M, 13B) models. The results, shown in Table 3, indicate that speculative decoding reduces the energy consumption and the time cost. This observation corroborates our claim to the energy efficiency of speculative decoding.
@@ 5 Experiments
Datasets and Models. We use five public datasets for evaluation: (1) ChatGPT-Prompt (Rashad, 2023), (2) ChatAlpaca (Bian et al., 2023), (3) CNN Dailymail (See et al., 2017), (4) Spider (Yu et al., 2018), and (5) MT-bench (Zheng et al., 2023). Table 8 in the Appendix shows more details of the datasets. Following previous studies (Kim et al., 2023), we use two public LLM families in our experiments: OPT (Zhang et al., 2022) and Llama-2 (Touvron et al., 2023). In this section, we set the large model to be OPT-13B and Llama-2-13B as they are the largest models that can run on a single 40GB GPU, and use Llama-68M (Miao et al., 2023) and OPT-125M as the small models. Appendix C reports additional experiment results with OPT-30B and Llama-2-chat-70B.
Baselines. For each pair of small and large models, we compare our method with four speculative decoding methods: vanilla speculative decoding (speculative(Lee et al., 2018; Chen et al., 2023), Spectr (Sun et al., 2023), SpecInfer (Miao et al., 2023), and BiLD (Kim et al., 2023). Our implementation of MTAD and all the baselines are based on a public implementation of speculative decoding (Bear, 2024). For each method, we let it generate at most 128 tokens for each input and run it for $1,000$ seconds. We open-sourced our code for reproduction. All the methods are stochastic with top-$k$ and top-$p$ sampling. The details of the hyper-parameters (e.g., $k$ and $p$) and machine configurations of the experiments can be found in the Appendix D, E, and F.
Appendix C reports additional experiments and ablation studies.
Table 4: Inference efficiency and output perplexity of different methods on ChatGPT-Prompt (CP), ChatAlpaca (CA), CNNDailyMail (CD), Spider (SP), and MT-Bench (MT) datasets. Bold numbers mark the best result, underlined numbers mark the second best.
@@
energy (J/token) $\downarrow$ 6.62$\pm$0.91 7.45$\pm$0.90 5.17$\pm$0.88 9.52$\pm$0.10 3.38$\pm$0.02
perplexity $\downarrow$ 3.64$\pm$0.11 3.15$\pm$0.06 3.64$\pm$0.08 3.64$\pm$0.11 2.06$\pm$0.06
OPT speed (token/s) $\uparrow$ 33.8$\pm$2.47 31.5$\pm$1.87 38.0$\pm$2.20 32.8$\pm$ 0.58 55.8$\pm$0.30
energy (J/token) $\downarrow$ 7.48$\pm$0.07 8.75$\pm$0.13 6.08$\pm$0.11 10.3$\pm$1.49 3.61$\pm$0.03
perplexity $\downarrow$ 5.47$\pm$0.11 4.51$\pm$0.09 5.27$\pm$0.09 5.12$\pm$0.01 3.00$\pm$0.09
CA Llama-2 speed (token/s) $\uparrow$ 31.6$\pm$0.35 28.8$\pm$0.20 27.7$\pm$0.29 26.5$\pm$0.49 44.1$\pm$0.25
energy (J/token) $\downarrow$ 6.98$\pm$0.15 7.99$\pm$0.15 7.20$\pm$0.08 7.52$\pm$0.32 4.72$\pm$0.03
perplexity $\downarrow$ 2.13$\pm$0.03 1.95$\pm$0.03 2.15$\pm$0.01 2.15$\pm$0.01 1.88$\pm$0.05
OPT speed (token/s) $\uparrow$ 35.6$\pm$0.45 38.5$\pm$0.93 28.4$\pm$0.34 31.4$\pm$0.39 49.6$\pm$0.42
energy (J/token) $\downarrow$ 5.74$\pm$0.11 5.12$\pm$0.06 6.24$\pm$0.11 8.68$\pm$1.83 4.03$\pm$0.02
perplexity $\downarrow$ 3.32$\pm$0.10 2.60$\pm$0.06 3.16$\pm$0.06 3.42$\pm$0.03 2.07$\pm$0.03
CD Llama-2 speed (token/s) $\uparrow$ 30.7$\pm$0.18 30.5$\pm$0.21 25.0$\pm$0.31 24.6$\pm$0.06 44.2$\pm$0.99
energy (J/token) $\downarrow$ 7.07$\pm$0.19 7.41$\pm$0.16 8.22$\pm$0.19 7.59$\pm$0.85 4.80$\pm$0.12
perplexity $\downarrow$ 2.87$\pm$0.08 2.93$\pm$0.03 3.06$\pm$0.11 2.92$\pm$0.09 2.63$\pm$0.10
OPT speed (token/s) $\uparrow$ 31.7$\pm$0.91 30.9$\pm$0.80 23.7$\pm$0.40 25.7$\pm$0.36 43.6$\pm$0.33
energy (J/token) $\downarrow$ 6.37$\pm$0.11 6.71$\pm$0.17 7.31$\pm$0.17 8.03$\pm$0.63 4.86$\pm$0.03
perplexity $\downarrow$ 3.97$\pm$0.06 3.74$\pm$0.09 4.04$\pm$0.07 3.92$\pm$ 0.34 3.17$\pm$0.06
SP Llama-2 speed (token/s) $\uparrow$ 24.0$\pm$0.28 26.2$\pm$0.08 24.2$\pm$0.29 23.8$\pm$0.20 26.4$\pm$0.28
energy (J/token) $\downarrow$ 10.75$\pm$0.02 9.84$\pm$0.07 11.0$\pm$0.08 11.0$\pm$0.76 9.01$\pm$0.07
perplexity $\downarrow$ 2.26$\pm$0.01 2.13$\pm$0.03 2.29$\pm$0.04 2.29$\pm$0.03 1.87$\pm$0.03
OPT speed (token/s) $\uparrow$ 24.6$\pm$0.30 29.9$\pm$0.55 19.8$\pm$0.13 24.1$\pm$0.10 34.4$\pm$0.46
energy (J/token) $\downarrow$ 15.6$\pm$3.55 13.6$\pm$3.07 20.1$\pm$2.52 16.9$\pm$2.75 11.7$\pm$2.36
perplexity $\downarrow$ 2.30$\pm$0.07 1.90$\pm$0.01 2.20$\pm$0.09 2.21$\pm$ 0.01 1.63$\pm$0.03
MT Llama-2 speed (token/s) $\uparrow$ 23.0$\pm$1.10 23.7$\pm$1.43 19.1$\pm$2.71 23.7$\pm$2.03 29.4$\pm$2.71
energy (J/token) $\downarrow$ 7.99$\pm$0.26 7.40$\pm$0.19 9.27$\pm$0.54 9.20$\pm$0.73 6.71$\pm$1.19
perplexity $\downarrow$ 3.64$\pm$0.51 3.44$\pm$0.76 3.64$\pm$0.51 3.63$\pm$0.50 2.21$\pm$0.18
OPT speed (token/s) $\uparrow$ 34.0$\pm$3.00 44.7$\pm$2.92 28.7$\pm$2.46 28.5$\pm$2.74 48.0$\pm$1.80
energy (J/token) $\downarrow$ 12.1$\pm$0.36 6.23$\pm$0.67 12.9$\pm$1.73 13.2$\pm$1.88 6.11$\pm$0.82
perplexity $\downarrow$ 2.02$\pm$0.40 1.50$\pm$0.27 1.97$\pm$0.38 1.99$\pm$ 0.33 1.10$\pm$0.03